OpenStudy (anonymous):

Write this equation in slope-intercept form,and then find the slope and y-intercept. 5y+3x=10

OpenStudy (anonymous):

5y+3x=10 5y=-3x+10 y=-3x/5+2 Can you tell what the slope and intercept are now?

OpenStudy (anonymous):

Yes,But I want to know how to find it..

OpenStudy (anonymous):

Find the equation? Just get the y by itself: that is, isolate it.

OpenStudy (anonymous):

So I divide 5y by 5 and 10 by 5 right? Then I'm left with y+3x=10 I think?